ID |
Date |
Author |
Subject |
Text |
 |
621
|
Sun Apr 28 22:52:06 2024 |
Marc | 0:00-08:00 Monday 29 April | Just another manic magic Monday,
....
Taking over Betool's AIDA shift. |
7x |
620
|
Sun Apr 28 20:35:57 2024 |
Betool Alayed | 16:0-00:00 28 April 2024 | 21:30 screenshots
and
23:24 screenshots |
10x |
619
|
Sun Apr 28 16:26:47 2024 |
Betool Alayed | 16:00-00:00 Sunday 28 April | 19:24 pm screenshots
and
17:18 pm screenshots |
10x |
618
|
Sun Apr 28 16:07:59 2024 |
TD | [How To] Recover DAQ when one FEE64 stops sending data - short(er) version | 0) Please infrom the 'on shift' crew that
you need to restart the AIDA DAQ
|
|
617
|
Sun Apr 28 15:49:58 2024 |
TD | DAQ restart short(er) cut |
16.45 aida04 stopped producing data
|
|
616
|
Sun Apr 28 15:09:34 2024 |
Betool Alayed | 16:0-00:00 28 April 2024 | 16:00 aida04 not producing data
power cycle all FEE64s to recover DAQ |
|
615
|
Sun Apr 28 06:57:59 2024 |
Magda Satrazani | 08:00-16:00 Sunday 28 April | Took over the shift from Marc.
|
36x |
614
|
Sat Apr 27 23:04:42 2024 |
Marc | 00:00-08:00 Sunday 28 April | Encore un matin, un matin pour ..... a
shift !!!ramping up again ! S4 temperature
not really going down either: 25C |
12x |
613
|
Sat Apr 27 15:13:06 2024 |
Norah | 16:00 - 00:00 Saturday 27 April | 14:50 Checks
Leakage current is 20.021
and HV status is Max Voltages ! Is it normal |
15x |
612
|
Sat Apr 27 04:59:31 2024 |
TD | 08:00-16:00 Saturday 27 April |
06.00 DAQ continues file S100_alpha/R20_740
FRS setting 162Eu
|
37x |
611
|
Fri Apr 26 23:04:51 2024 |
Marc | 00:00-08:00 Saturday 27 April | It's a new day, it's a new shift
and I'm feeeeeliiiiing goooood !!!!!
0.15 Checks:
DSSSD bias |
12x |
610
|
Fri Apr 26 15:34:00 2024 |
RDP | 16:00 - 00:00 Friday 26 April | This appeared in the system log:
Apr 26 16:30:51 aidas-gsi smartd[1076]:
Device: /dev/sda [SAT], 1 Currently unreadable |
6x |
609
|
Fri Apr 26 08:15:59 2024 |
TD | [How To] Recover DAQ when one FEE64 stops sending data |
Occasionally, one (sometimes more) FEE64(s)
will stop producing data whilst the DAQ is |
|
608
|
Fri Apr 26 07:01:26 2024 |
Norah | 08:00 - 16:00 Friday 26 April | 10:00 Checks:
All seems smooth.
DSSD bias and leakage current |
    |
607
|
Thu Apr 25 22:45:30 2024 |
Tom, Muneerah, Norah | Over night shift AIDA 26 Fri |
01:02
Aida04 was zero not reading had |
9x |
606
|
Thu Apr 25 15:02:50 2024 |
Marc | 16:00 - 0:00 Thursday 25 April |
16:20:
Spill length:4s (was 3.5s
before) - Extraction time 1.5s (see attachement
#1)
|
8x |
605
|
Thu Apr 25 08:26:50 2024 |
TD | 08:00-16:00 Thursday 25 April | 09.56 DAQ continues data file S100_alpha/R18_387
FRS switching settings
|
42x |
604
|
Wed Apr 24 21:39:51 2024 |
TD | 00:00-08:00 Thursday 25 April | 22.43 24.4.24
DSSSD bias & leakage current OK - attachment |
12x |
603
|
Wed Apr 24 18:44:05 2024 |
PP | 19:30 checks | All seems smooth.
Screenshots attached. |
8x |
602
|
Wed Apr 24 15:37:59 2024 |
PP | 16:00-00:00 shift Wednesday 24 April | All seems OK.
Screenshots attached. |
8x |